Q5. Katherine advertise to sell cookies for $4 a dozen. She sells 50 dozens and decided that she can charge more. She raise the price to $6 a dozen and sell 40 dozen. What is elasticity of demand? Assuming that the elasticity of demand is constant. How many would she sell if the price were $10 a box?


1
Expert's answer
2021-04-05T07:23:30-0400

1 dozen = $4

50 dozens = $4"\u00d7" 50 = $200


at the price of $6 per dozen,

40 dozens = $6"\u00d7" 40 = $240


"Q_{o}=" old quantity(50); "Q_{n}=" new quantity(40); "P_{o}=" old price(200); "P_{n}=" new price(240)


"\epsilon=\frac{\frac{Q_{n}-Q_{o}}{Q_{o}}×100}{\frac{P_{n}-P_{o}}{P_{o}}×100}"


"\epsilon=\frac{\frac{40-50}{50}×100}{\frac{240-200}{200}×100}"

"\epsilon=" 1 (elasticity is unitary)


if elasticity is constant, i.e. "\epsilon=1", then the new quantity to be sold at $10 a box will be "x" and the new price, "10x"

assume that

"\frac{\frac{Q_{n}-Q_{o}}{Q_{o}}×100}{\frac{P_{n}-P_{o}}{P_{o}}×100}=1"


"\frac{\frac{40-x}{40}×100}{\frac{240-10x}{240}×100}=1"

"x" , (quantity to be sold at price $10 per box) is 22.2 dozens